You have a coin and a 6 sided die. A. What is the probability of getting a heads on a coin flip? B. What is the probability of getting a 4 on a diet Roll? C. What is the probability of getting tails on a coin flip and a 3 on a diet roll? D. What is the probability of getting a 2 or 4 or 5 on a diet ROLL? E. What is the probability of rolling a die three times and only getting a 4 on the third ROLL?

Found 2 solutions by Boreal, math_helper:
Answer by Boreal(15235)

The first is 1/2
A 4 appears with 1/6 probability
tails and a 3 is (1/2)(1/6)=1/12, since they are independent.
a 2,4,5 are 3 of 6 possibilities so probability is 1/2
finally getting a 4 is (5/6)(5/6)(1/6)=25/216 probability, since they are independent. What you don't want *2, multiplied by what you do want.

Answer by math_helper(2461)


A. P(heads flipped) = 1/2
B. P(4 rolled) = 1/6
C. P(Tails flipped AND 3 rolled on die) = (1/2)*(1/6) = 1/12
D. P(2,4,or 5 rolled on die) = 3/6 = 1/2
E. P(die rolled three times and a 4 shows on 3rd roll only) = (5/6)*(5/6)*(1/6) = 25/216